OpenStudy (anonymous):

how to find the highest point on the cardioid r = 1 + cos θ.

OpenStudy (anonymous):

when you say the highest point, you mean, maximum "y" value?

OpenStudy (anonymous):

yeah i thnk soo

OpenStudy (anonymous):

or the maximum "r" = distance of outermost point from the center?

OpenStudy (anonymous):

not sure bt da answer shud cme to (3/2, π/3)...dnt knw how??

OpenStudy (anonymous):

ok \[r(\theta)=1+\cos\theta\\y(\theta)=r(\theta)\cdot\sin\theta\\\implies y(\theta)=\sin\theta+\cos\theta\sin\theta\] we are maximizing "y" with theta. os, its first derivative is "0"

OpenStudy (anonymous):

\[y'=\cos\theta+\cos^2\theta-\sin^2\theta=0\\\cos\theta+\cos^2\theta-(1-\cos^2\theta)=0\] solve for theta

OpenStudy (anonymous):

ohhhk.. thanks alot...

OpenStudy (anonymous):

yw